# Imaging the positron annihilation line with 20-year INTEGRAL/SPI observations

### v2

@@ Abstract
Aims. We analyze 20 years of INTEGRAL/SPI observations to produce the most sensitive all-sky map of the 511 keV line emission to date, aiming to reveal new features and provide refined measurements of known sources.
Methods. We perform image deconvolution using the Richardson-Lucy algorithm and employ a bootstrap analysis to evaluate statistical uncertainties of fluxes from regions of interest. Systematic uncertainties in the parameter choices for the image reconstruction are also considered. We utilize GPU acceleration to enable this computationally intensive data analysis.
Results. The reconstructed image successfully recovers the basic morphological features reported in model-fitting studies: a bright central component, a broad bulge, and an elongated disk component along the Galactic plane. We also report hints of new spatial features in the reconstructed image, including an asymmetric structure in the broad bulge emission and 511 keV emission potentially associated with massive stars from the Scorpius–Centaurus and other OB associations. While the significance of these new features is marginal ($\sim 2\sigma$), they are spatially consistent with 26Al emission from massive stars in that region, suggesting that this 511 keV emission originates from its $\beta^{+}$ decay.
Conclusions. Our 20-year dataset provides the most detailed 511 keV emission map to date, reproducing the global structures suggested in the model-fitting approach while revealing hints of new spatial features. These findings provide new insights into the origin of Galactic positrons and the propagation of low-energy positrons in the interstellar medium. Future MeV gamma-ray observations, such as with COSI, are expected to confirm the reported features and shed further light on the nature of positrons in our Galaxy.
@@ 1 Introduction
Positrons, the antimatter counterparts of electrons, play fundamental roles in high-energy astrophysics. Their existence was first predicted theoretically by Dirac (1931) and experimentally discovered by Anderson (1932). When a positron encounters an electron, they can either directly annihilate to produce two photons with 511 keV energy each if the particles are sufficiently cooled, or form a bound state called positronium (Ps). Ps eventually decays, producing two 511 keV photons (para-Ps) or three photons with a continuum spectrum below 511 keV (ortho-Ps), depending on its spin states. This characteristic spectral signature provides a unique probe of antimatter in the Universe. Understanding the production, propagation, and annihilation of positrons is crucial for studying various astrophysical phenomena, from stellar evolution to particle astrophysics.
@@
The discovery of the Galactic 511 keV emission dates back to balloon experiments (e.g., Haymes et al. 1969; Johnson & Haymes 1973). Since then, it has been measured by multiple satellite missions for over 50 years. OSSE aboard the Compton Gamma Ray Observatory (Johnson et al. 1993) provided the first maps of the 511 keV emissions, revealing its diffuse nature at the Galactic Center (e.g., Purcell et al. 1997). After that, INTEGRAL/SPI (Winkler et al. 2003; Vedrenne et al. 2003) revolutionized our understanding with its high energy and angular resolution by unveiling its complex morphology with increasing exposure time since its launch in 2002.
@@
Several key observational features have been obtained from these studies (e.g., Churazov et al. 2020; Prantzos et al. 2011). Around the Galactic center, the 511 keV emission is dominated by a diffuse bulge component with an extent of 3-12 degrees, potentially accompanied by a central point source (Skinner et al. 2015). INTEGRAL/SPI has also detected extended disk emission along the Galactic plane. The latest measurements (Skinner et al. 2015; Siegert et al. 2016) fitted data with analytical spatial models for these components, and reported the fluxes of the bulge, central point source, and the disk to be $(8.9-10.1)\times 10^{-4}$, $(0.6-1.2)\times 10^{-4}$, and $(13.1-20.1)\times 10^{-4}\leavevmode\nobreak\ \mathrm{ph\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$, respectively. The ratio of the 511 keV emission from the bulge is comparable to that from the disk emission. Such a large bulge-to-disk ratio is a unique feature not seen in other wavelengths. Note that Skinner et al. (2015) claimed that the Galactic disk emission has a scale height of 3, which is the $\sigma$ parameter of the two-dimensional Gaussian function for the disk component, vertical to the Galactic plane. In contrast, Siegert et al. (2016) reported that it has a larger scale height of 10.5. We refer to the former and latter as the thin and thick disk models, respectively. Such a difference is probably caused by different data and background treatments.
@@
Despite decades of study, the origin of the Galactic positrons remains one of the long-term mysteries in high-energy astrophysics, the so-called “positron puzzle” (Siegert 2023). Various sources have been proposed to explain the Galactic 511 keV emission, each with a different spatial distribution, e.g., $\beta^{+}$ decay of radioisotopes emitted from massive stars (Diehl et al. 2006, 2010; Pleintinger et al. 2023), core-collapse supernovae (Siegert et al. 2015), type Ia supernovae (Milne et al. 1999; Prantzos 2006; Siegert 2017), and classical novae (Leung & Siegert 2022); e-e+ production in X-ray binaries and pulsars (Guessoum et al. 2006; Weidenspointner et al. 2008; Wang et al. 2006); e-e+ production near the hot inner disk of Sgr A (for the bulge emission, Totani 2006); light dark matter models as an exotic scenario (Finkbeiner & Weiner 2007; Siegert et al. 2022; Sheng et al. 2024; Watanabe et al. 2025). Through cosmic-ray interactions, positrons can also be produced by nearby sources like the Moon and asteroids (Moskalenko et al. 2008; Siegert 2024; Fujiwara et al. 2025). However, no single scenario fully explains the observed flux and spatial distributions, suggesting the 511 keV line likely originates from a combination of (some of) these sources.
@@
In this work, we analyze the spatial features using 20 years of INTEGRAL/SPI observations of the 511 keV emission line, representing the highest available photon statistics. We employ the Richardson-Lucy algorithm for image reconstruction, which is particularly suited for recovering the 511 keV map from the complex data and response of the coded-masked telescope of SPI. Additionally, we also optimize the background components simultaneously during the image reconstruction process. Furthermore, we implement bootstrap methods to quantify the significance of detected spatial structures and their flux uncertainties. Such an analysis was not previously performed due to the heavy computation required. Here, we use GPU-accelerated computing to achieve reasonable uncertainty estimates.
The paper is organized as follows. Section 2 describes the SPI observations and data reduction. Section 3 presents our image deconvolution methods, background modeling, and uncertainty evaluations. Section 4 presents the updated 511 keV map, revealing potential new spatial features. Section 5 discusses these findings with possible interpretations and future prospects with upcoming MeV gamma-ray telescopes. Finally, we conclude our work in Section 6.
@@ 2 Observations and data reduction
The INTErnational Gamma-Ray Astrophysics Laboratory (INTEGRAL Winkler et al. 2003) has been observing the gamma-ray sky since October 17, 2002, and ceased its scientific operations on February 28, 2025. The spectrometer SPI (Vedrenne et al. 2003) onboard the satellite was the most sensitive gamma-ray spectrometer telescope with a coded aperture mask ever built. It measured photons in the energy range from 18 keV to beyond 8 MeV with an unprecedented energy resolution of 0.41% (FWHM) at 511 keV and an angular resolution of $2.7^{\circ}$ (FWHM) (Diehl et al. 2018). Due to cosmic-ray bombardment, about 99% of the recorded events are typically due to instrumental background. Additionally, this radiation dosage from cosmic rays and trapped particles in the Van Allen radiation belts degrades the energy resolution of the 19 SPI’s germanium detectors. With annealing procedures, heating up the detectors to above $100^{\circ}\mathrm{C}$ about every half year, the resolution is restored almost completely (Roques et al. 2003), except for a long-term drift that worsens the resolution over the 22 mission years by about 10% to 0.46% at 511 keV.
In this work, we analyze 20 years of INTEGRAL/SPI data covering the orbital revolutions 43–2679. It spans from February 19, 2003 to August 29, 2023 (ISDC Julian Date 1145–8641). Using the ISDC standard tools within its OSA package [^1] https://www.astro.unige.ch/integral/analysis#Software, we select pointings according to several selection criteria: We require that each pointing is at least 120 s long. Further, five hours before and after a solar flare are excluded, and individually high count rates, due to flaring sources or particle events in the detectors, are also excluded. More details are described in Appendix A. This leaves us with 143,833 pointings. Using a first-order maximum likelihood fit using a background model and a sky model, we find that 206 individual pointings show an erratic behavior in that the background plus sky combination does not describe the data properly, showing residuals with $\pm 10\sigma$ or above. We filter out these pointings and perform an additional maximum likelihood with the same background and sky models, which results in a screened dataset. After the data screening, we obtained 143,627 pointing observations with a total gross exposure time of 367.4 Ms. The typical exposure per observation ranges from $\sim$2 ks to $\sim$4 ks, with some long staring observations exceeding 20 ks. It yields an average exposure of 2.56 ks per observation.
@@
With such a long exposure, continuum emissions from bright point sources can contaminate our analysis even within a narrow energy band for this gamma-ray line analysis. We account for this by modeling and subtracting these sources as “background” components. The exposure is not uniformly distributed in high latitude regions, e.g., $|b|>40^{\circ}$, which may cause artifacts in the image reconstruction discussed later. We also want to mention that upcoming MeV satellites such as COSI (Tomsick et al. 2023) will address this issue through all-sky surveys with a wide field-of-view (FoV) Compton telescope.
@@
Our analysis uses only single-detector events (SE) because they dominate the 511 keV signal data, and the contribution from multiple-detector events (ME) can be negligible. Note that SE are the events that deposit their energy in a single detector, while ME are those that deposit their energy in more than one detector. More details are found in Vedrenne et al. (2003). The data have been analyzed by sorting the events in a 3-dimensional data space, spanned by the (calibrated) event energy, the detector number, and the SPI pointing number (see Sect. 3.1). An initial energy binning of 0.5 keV is used by the standard pipeline from ISDC, well below the instrumental energy resolution. Our final imaging analysis is carried out in a single 6 keV wide energy bin from 508–514 keV. For the analysis of point sources and their contribution to the 511 keV bin, we use a spectral range from 400–640 keV with 40 energy bins of 6 keV width each. In this way, we can estimate the continuum flux of different sources (see Sect. 3.1.2).
@@ 3.1.1 Response
Next, we calculate the total expected counts in each detector. The response of the coded mask aperture is modeled as the matrix $R^{\mathrm{mask}}_{k,d}$, representing the effective area of detector $d$ for a source at pixel $k$. With the livetime $T_{p,d}$ of each detector and each pointing, the total expected counts can be calculated as $T_{p,d}\sum_{k}\sum_{j}R^{\mathrm{mask}}_{k,d}C_{p,j,k}\lambda_{j}$. Finally, we consider the response of the germanium detectors. The energy redistribution matrix $R^{\mathrm{rmf}}_{e}$ represents the probability that a detected 511 keV gamma ray is measured in energy bin $e$. Consequently, the expected count $\epsilon^{\mathrm{source}}_{p,d,e}$ at each pointing, detector, and energy bin from a given 511 keV map can be calculated as:
@@
$=$ $\sum_{j}R_{p,d,e,j}\lambda_{j}\leavevmode\nobreak\ ,$ (2)
where for simplicity we define $R_{p,d,e,j}$ as:
$R_{p,d,e,j}\equiv T_{p,d}\sum_{k}R^{\mathrm{rmf}}_{e}R^{\mathrm{mask}}_{k,d}C_{p,j,k}\leavevmode\nobreak\ .$ (3)
@@ 3.1.2 Background modeling
The instrumental background in SPI is the most dominant component at all energies except for gamma-ray bursts or bright transients. We use the method by Siegert et al. (2019), who define the instrumental response and background database (Diehl et al. 2018), which is still maintained until the end of the INTEGRAL mission. This database uses the raw count spectra per detector and revolution in 0.5 keV binning and fits all the spectral features using about 600 gamma-ray lines on top of a multiple broken power-law continuum from 20 keV to 2 MeV. At SPI’s high energy range (2–8 MeV), there are about 1000 more gamma-ray lines to be evaluated, but which cannot be done with high accuracy on the orbital time scale of three days. Around the 511 keV line, and in particular for the energy bin of interest from 508–514 keV, there are two gamma-ray lines on top of a continuum. Since the raw spectra are dominated by instrumental background, it was found that the relative count rate of different background components in each detector stays constant over time. These “detector patterns” are typically combined to contributions from instrumental lines ($B^{\mathrm{line}}_{p,d,e}$) and instrumental continuum ($B^{\mathrm{cont}}_{p,d,e}$) to avoid over-fitting.
@@
For the potentially bright point sources, we fit the data from 400 to 640 keV with power-law functions and estimate the integrated flux from 508–514 keV. The point sources are included as background components to only show the truly diffuse emission, and are listed with the estimated fluxes in Table 3. Although some sources are known as time-variable sources, considering their time-variability in our analysis did not affect the results. Thus, we used the time-averaged flux in this background modeling.
@@
$\epsilon^{\mathrm{bkg}}_{p,d,e}=b^{\mathrm{line}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{line}}_{p,d,e}+b^{\mathrm{cont}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{cont}}_{p,d,e}+\sum_{m}B^{\mathrm{source},m}_{p,d,e}\leavevmode\nobreak\ .$ (4)
@@
Table 1: List of bright sources subtracted in the image reconstruction.
@@
 [^3] The fluxes shown here are integrated over the energy range of 508 keV to 514 keV. The source locations are given in Galactic coordinates.

### v3

# Imaging the positron annihilation line with 20 years of INTEGRAL/SPI observations

@@ Abstract
Aims. We analyzed 20 years of INTEGRAL/SPI observations to produce the most sensitive all-sky map of the 511 keV line emission to date with the aim of revealing new features and providing refined measurements of known sources.
Methods. We performed image deconvolution using the Richardson-Lucy algorithm and employed a bootstrap analysis to evaluate statistical uncertainties of fluxes from regions of interest. Systematic uncertainties in the parameter choices for the image reconstruction were also considered. We utilized GPU acceleration to enable this computationally intensive data analysis.
Results. The reconstructed image successfully recovers the basic morphological features reported in model-fitting studies: a bright central component, a broad bulge, and an elongated disk component along the Galactic plane. We also report hints of new spatial features in the reconstructed image, including an asymmetric structure in the broad bulge emission and a 511 keV emission potentially associated with massive stars from the Scorpius–Centaurus and other OB associations. While the significance of these new features is marginal ($\sim 2\sigma$), they are spatially consistent with 26Al emission from massive stars in that region, suggesting that this 511 keV emission originates from its $\beta^{+}$ decay.
Conclusions. Our 20-year dataset provides the most detailed 511 keV emission map to date, reproducing the global structures suggested in the model-fitting approach while also revealing hints of new spatial features. These findings provide new insights into the origin of Galactic positrons and the propagation of low-energy positrons in the interstellar medium. Future MeV gamma-ray observations, such as with COSI, are expected to confirm the reported features and shed further light on the nature of positrons in our Galaxy.
@@ 1 Introduction
Positrons, the antimatter counterparts of electrons, play fundamental roles in high-energy astrophysics. Their existence was first predicted theoretically by Dirac (1931) and experimentally discovered by Anderson (1932). When a positron encounters an electron, they can either directly annihilate to produce two photons with 511 keV energy each if the particles are sufficiently cooled or form a bound state called positronium (Ps). Ps eventually decays, producing two 511 keV photons (para-Ps) or three photons with a continuum spectrum below 511 keV (ortho-Ps), depending on its spin states. This characteristic spectral signature provides a unique probe of antimatter in the Universe. Understanding the production, propagation, and annihilation of positrons is crucial for studying various astrophysical phenomena, from stellar evolution to particle astrophysics.
@@
The discovery of the Galactic 511 keV emission dates back to balloon experiments (e.g., Haymes et al. 1969; Johnson & Haymes 1973). Since then, it has been measured by multiple satellite missions for over 50 years. OSSE aboard the Compton Gamma Ray Observatory (Johnson et al. 1993) provided the first maps of the 511 keV emission, revealing its diffuse nature at the Galactic Center (e.g., Purcell et al. 1997). After that, INTEGRAL/SPI (Winkler et al. 2003; Vedrenne et al. 2003) revolutionized our understanding with its high energy and angular resolution by unveiling its complex morphology with increasing exposure time since its launch in 2002.
@@
Several key observational features have been obtained from these studies (e.g., Churazov et al. 2020; Prantzos et al. 2011). Around the Galactic center, the 511 keV emission is dominated by a diffuse bulge component with an extent of 3-12 degrees and is potentially accompanied by a central point source (Skinner et al. 2015). INTEGRAL/SPI has also detected extended disk emission along the Galactic plane. The latest measurements (Skinner et al. 2015; Siegert et al. 2016) fitted data with analytical spatial models for these components and reported the fluxes of the bulge, central point source, and the disk to be $(8.9-10.1)\times 10^{-4}$, $(0.6-1.2)\times 10^{-4}$, and $(13.1-20.1)\times 10^{-4}\penalty 10000\ \mathrm{ph\penalty 10000\ cm^{-2}\penalty 10000\ s^{-1}}$, respectively. The ratio of the 511 keV emission from the bulge is comparable to that from the disk emission. Such a large bulge-to-disk ratio is a unique feature not seen in other wavelengths. We note that Skinner et al. (2015) claimed that the Galactic disk emission has a scale height of 3, which is the $\sigma$ parameter of the two-dimensional Gaussian function for the disk component vertical to the Galactic plane. In contrast, Siegert et al. (2016) reported that it has a larger scale height of 10.5. We refer to the former and latter as the thin and thick disk models, respectively. Such a difference is probably caused by different data and background treatments.
@@
Despite decades of study, the origin of the Galactic positrons remains one of the long-term mysteries in high-energy astrophysics, the so-called positron puzzle (Siegert 2023). Various sources have been proposed to explain the Galactic 511 keV emission, each with a different spatial distribution, for example, $\beta^{+}$ decay of radioisotopes emitted from massive stars (Diehl et al. 2006, 2010; Pleintinger et al. 2023), core-collapse supernovae (Siegert et al. 2015), type Ia supernovae (Milne et al. 1999; Prantzos 2006; Siegert 2017), and classical novae (Leung & Siegert 2022); e-e+ production in X-ray binaries and pulsars (Guessoum et al. 2006; Weidenspointner et al. 2008; Wang et al. 2006); e-e+ production near the hot inner disk of Sgr A (for the bulge emission, Totani 2006); light dark matter models as an exotic scenario (Finkbeiner & Weiner 2007; Siegert et al. 2022; Sheng et al. 2024; Watanabe et al. 2025). Through cosmic-ray interactions, positrons can also be produced by nearby sources such as the Moon and asteroids (Moskalenko et al. 2008; Siegert 2024; Fujiwara et al. 2025). However, no single scenario fully explains the observed flux and spatial distributions, suggesting the 511 keV line likely originates from a combination of (some of) these sources.
@@
In this work, we analyze the spatial features using 20 years of INTEGRAL/SPI observations of the 511 keV emission line, representing the highest available photon statistics. We employed the Richardson-Lucy algorithm for image reconstruction, which is particularly suited for recovering the 511 keV map from the complex data and response of the coded-masked telescope of SPI. Additionally, we optimized the background components simultaneously during the image reconstruction process. Furthermore, we implemented bootstrap methods to quantify the significance of detected spatial structures and their flux uncertainties. Such an analysis has not been performed previously due to the heavy computation required. Here, we use GPU-accelerated computing to achieve reasonable uncertainty estimates.
The paper is organized as follows. In Sect. 2, we describe the SPI observations and data reduction. Section 3 presents our image deconvolution methods, background modeling, and uncertainty evaluations. Section 4 presents the updated 511 keV map, revealing potential new spatial features. Section 5 discusses these findings with possible interpretations and future prospects with upcoming MeV gamma-ray telescopes. Finally, we conclude our work in Sect. 6.
@@ 2 Observations and data reduction
The INTErnational Gamma-Ray Astrophysics Laboratory (INTEGRAL Winkler et al. 2003) has been observing the gamma-ray sky since October 17, 2002, and ceased its scientific operations on February 28, 2025. The spectrometer SPI (Vedrenne et al. 2003) on board the satellite was the most sensitive gamma-ray spectrometer telescope with a coded aperture mask ever built. It measured photons in the energy range from 18 keV to beyond 8 MeV with an unprecedented energy resolution of 0.41% (full width at half maximum; FWHM) at 511 keV and an angular resolution of $2.7^{\circ}$ (FWHM) (Diehl et al. 2018). Due to cosmic-ray bombardment, about 99% of the recorded events are typically due to instrumental background. Additionally, this radiation dosage from cosmic rays and trapped particles in the Van Allen radiation belts degrades the energy resolution of the 19 SPI’s germanium detectors. With annealing procedures, heating up the detectors to above $100^{\circ}\mathrm{C}$ about every half year restored the resolution almost completely (Roques et al. 2003), except for a long-term drift that worsened the resolution over the 22 mission years by about 10% to 0.46% at 511 keV.
In this work, we analyze 20 years of INTEGRAL/SPI data covering the orbital revolutions 43–2679. It spans from February 19, 2003, to August 29, 2023 (ISDC Julian Date 1145–8641). Using the ISDC standard tools within its OSA package, [^1] https://www.astro.unige.ch/integral/analysis#Software we selected pointings according to several selection criteria: We required that each pointing is at least 120 s long. Further, we excluded five hours before and after solar flares, and individual high count rates due to flaring sources or particle events in the detectors. More details are described in Appendix A. This left us with 143,833 pointings. Using a first-order maximum likelihood fit using a background model and a sky model, we found that 206 individual pointings show an erratic behavior in that the background plus sky combination does not describe the data properly, showing residuals with $\pm 10\sigma$ or above. We filtered out these pointings and performed an additional maximum likelihood with the same background and sky models, which resulted in a screened dataset. After the data screening, we obtained 143,627 pointing observations with a total gross exposure time of 367.4 Ms. The typical exposure per observation ranges from $\sim$2 ks to $\sim$4 ks, with some long staring observations exceeding 20 ks. The screened dataset yields an average exposure of 2.56 ks per observation.
@@
With such a long exposure, continuum emissions from bright point sources can contaminate our analysis even within a narrow energy band for this gamma-ray line analysis. We account for this by modeling and subtracting these sources as “background” components. The exposure is not uniformly distributed in high latitude regions, for example, $|b|>40^{\circ}$, which may cause artifacts in the image reconstruction discussed later. We also want to mention that upcoming MeV satellites such as COSI (Tomsick et al. 2023) will address this issue through all-sky surveys with a wide field-of-view (FoV) Compton telescope.
@@
Our analysis uses only single-detector events (SEs) because they dominate the 511 keV signal data, and the contribution from multiple-detector events (MEs) can be negligible. We note that SEs are the events that deposit their energy in a single detector, while MEs are those that deposit their energy in more than one detector. More details are found in Vedrenne et al. (2003). The data were analyzed by sorting the events in a three-dimensional data space, spanned by the (calibrated) event energy, the detector number, and the SPI pointing number (see Sect. 3.1). An initial energy binning of 0.5 keV was used by the standard pipeline from ISDC, well below the instrumental energy resolution. Our final imaging analysis was carried out in a single 6 keV wide energy bin from 508–514 keV. For the analysis of point sources and their contribution to the 511 keV bin, we used a spectral range from 400–640 keV with 40 energy bins of 6 keV width each. In this way, we estimated the continuum flux of different sources (see Sect. 3.1.2).
@@ 3.1.1 Response
Next, we calculate the total expected counts in each detector. The response of the coded mask aperture is modeled as the matrix $R^{\mathrm{mask}}_{k,d}$, representing the effective area of detector $d$ for a source at pixel $k$. With the livetime $T_{p,d}$ of each detector and each pointing, the total expected counts can be calculated as $T_{p,d}\sum_{k}\sum_{j}R^{\mathrm{mask}}_{k,d}C_{p,j,k}\lambda_{j}$. Finally, we consider the response of the germanium detectors. The energy redistribution matrix $R^{\mathrm{rmf}}_{e}$ represents the probability that a detected 511 keV gamma ray is measured in energy bin $e$. Consequently, the expected count $\epsilon^{\mathrm{source}}_{p,d,e}$ at each pointing, detector, and energy bin from a given 511 keV map can be calculated as
@@
$=$ $\sum_{j}R_{p,d,e,j}\lambda_{j}\penalty 10000\ ,$ (2)
where for simplicity we define $R_{p,d,e,j}$ as
$R_{p,d,e,j}\equiv T_{p,d}\sum_{k}R^{\mathrm{rmf}}_{e}R^{\mathrm{mask}}_{k,d}C_{p,j,k}\penalty 10000\ .$ (3)
@@ 3.1.2 Background modeling
The instrumental background in SPI is the most dominant component at all energies except for gamma-ray bursts or bright transients. We used the method by Siegert et al. (2019), who defined the instrumental response and background database (Diehl et al. 2018), which was maintained until the end of the INTEGRAL mission. This database uses the raw count spectra per detector and revolution in 0.5 keV binning and fits all the spectral features using about 600 gamma-ray lines on top of a multiple broken power-law continuum from 20 keV to 2 MeV. At SPI’s high energy range (2–8 MeV), there are about 1000 more gamma-ray lines to be evaluated, but which cannot be done with high accuracy on the orbital time scale of three days. Around the 511 keV line, and in particular for the energy bin of interest from 508–514 keV, there are two gamma-ray lines on top of a continuum. Since the raw spectra are dominated by instrumental background, it was found that the relative count rate of different background components in each detector stays constant over time. These “detector patterns” are typically combined to contributions from instrumental lines ($B^{\mathrm{line}}_{p,d,e}$) and instrumental continuum ($B^{\mathrm{cont}}_{p,d,e}$) to avoid over-fitting.
@@
For the potentially bright point sources, we fitted the data from 400 to 640 keV with power-law functions and estimated the integrated flux from 508–514 keV. The point sources are included as background components to only show the truly diffuse emission, and are listed with the estimated fluxes in Table 3. Although some sources are known as time-variable sources, considering their time-variability in our analysis did not affect the results. Thus, we used the time-averaged flux in this background modeling.
@@
$\epsilon^{\mathrm{bkg}}_{p,d,e}=b^{\mathrm{line}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{line}}_{p,d,e}+b^{\mathrm{cont}}_{\mathrm{rev}(p)}\sum_{p,d,e}B^{\mathrm{cont}}_{p,d,e}+\sum_{m}B^{\mathrm{source},m}_{p,d,e}\penalty 10000\ .$ (4)
@@
Table 1: List of bright sources subtracted in the image reconstruction. [^3] The fluxes shown here are integrated over the energy range of 508 keV to 514 keV. The source locations are given in Galactic coordinates.